# Context-free graphs and their transition groups

### v2

@@ 4. Transition groups of inverse graphs and co-context-free groups
Let us prove the characterization. It is straightforward to check that the transition group of the Cayley graph $\operatorname{Cay}(G;A)$ is isomorphic to $G$. Thus, if there is a covering $\psi:\operatorname{Cay}(G;A)\to\Gamma$ onto a context-free graph $\Gamma$, then by the first part of this proposition $\mathcal{G}(\Gamma)\simeq G$. Conversely, if $G=\mathcal{G}(\Gamma)$ and we fix a root $x_{0}$ in $\Gamma$, then the morphism $\psi:\operatorname{Cay}(G;A)\to\Gamma$ defined on the vertices by $\psi(v)=y$ whenever $1\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}v$ is a walk in $\operatorname{Cay}(G;A)$ and $y=x_{0}\cdot g$ and extended to a morphism of inverse $A$-digraphs, is a covering that is also transition-preserving. ∎
This class is closed under taking finitely generated subgroups, direct products and finite extensions. In order to prove that, we need some preliminary lemma.
We fix $H$ a subgroup of $G$ that is finitely generated by $B\subseteq A^{*}$ and we set $n:=\max_{B}\ell(b)$ with $\ell(b)$ the length in terms of $A$.
Lemma 4.13.
Let $g\in H$ be a non-trivial element. Then, there exist an end-cone $\Gamma(v,x_{0})$ and a vertex $\widetilde{v}\in\Gamma(v,x_{0})$ with $d(v,\widetilde{v})\leq n$ such that the infinite walk $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(v,x_{0})$ for some $\mathring{g}\in S_{B}(\{g\})$ and $\widetilde{v}\cdot\mathring{g}\neq\widetilde{v}$.
Proof.
By Lemma 4.5 there exists an end-cone $\Gamma(v,x_{0})$ such that the infinite walk $\pi=x\smash{\stackrel{{\scriptstyle\text{\tiny{$\widehat{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(v,x_{0})$. Now, by definition of $n$, up to removing a walk of at most $n$ edges to $\pi$, we obtain a walk $v\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$ with $d(v,\widetilde{v})\leq n$ and $\mathring{g}\in S_{B}(\{g\})$. ∎
Proposition 4.14.
The class of CF-TR is closed by taking finitely generated subgroups.
Proof.
Suppose first that we are dealing with one complete context-free graph, so that $H$ is a finitely generated subgroup of $\mathcal{G}(\Gamma)$, where $\Gamma$ is an inverse graph that is also context-free. Let $B\subseteq A^{*}$ be the involutive closed generating set and let $\Delta_{1},\ldots,\Delta_{N}$ be the frontier points of the end-cone types $\{\Gamma_{1},\ldots,\Gamma_{N}\}$. We set $\Delta_{i}^{n}:=\{v\in\Gamma_{i}\mid d(x,v)\leq n\mbox{ for some }x\in\Delta_{i}\}$. With a slight abuse of notation we identify the frontier vertices of $\Delta_{i}$ with the frontier vertices of an end cone $\Gamma(v,x_{0})$ of $\Gamma$ that is end-isomorphic to $\Gamma_{i}$. We do the same for the set $\Delta_{i}^{n}$. By definition of transition group, we need to show that $H$ is the transition group of a finite union of complete context-free $B$-digraphs $\Theta_{B}$. We claim
$\Theta_{B}=\bigcup_{i=1}^{N}\bigcup_{v\in\Delta_{i}^{n}}(\Gamma^{B},v)$
where $(\Gamma^{B},v)$ is given by the vertices $y$ of $\Gamma$ such that there exists $w\in B^{*}$ and $v\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}y$. The edges are $y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}y_{2}$ whenever $w\in B$ and there exists a walk labeled by $w$ from $y_{1}$ to $y_{2}$ in $\Gamma$. It is not difficult to see that since $B$ is involutive closed and $\Gamma$ is inverse, then also $(\Gamma^{B},v)$ is inverse. Moreover, it is complete and context-free.
Let us show that the complement of the word problem of $H$ and $\mathcal{G}(\Theta_{B})$ coincide, as this will of course imply that their word problems coincide. Let $g\in G(\Theta_{B})$ be a non trivial element, then there exists $\widetilde{v}\in(\Gamma^{B},v)$ for some $v\in\Delta_{i}^{n}$ such that $\widetilde{v}\cdot g\neq\widetilde{v}$. It follows that $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime}$ with $\widetilde{v}\neq v^{\prime}$ is also a walk in $\Gamma$ and so $g\neq\mathds{1}_{H}$ in $H$.
Conversely, suppose that $\mathds{1}_{H}\neq g\in H$, by Lemma 4.13 there exists an end-cone $\Gamma(x,x_{0})$ and a vertex $v\in\Gamma(x,x_{0})$ with $d(x,v)\leq n$ such that $v\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(x,x_{0})$ and $v\cdot\mathring{g}\neq v$ for some $\mathring{g}\in S_{B}(\{g\})$. Now $\Gamma(x,x_{0})$ is end-isomorphic to some $\Gamma_{i}$, hence we can find $\widetilde{v}\in\Delta_{i}^{n}$ such that $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma_{i}$ and $\widetilde{v}\cdot\mathring{g}\neq\widetilde{v}$. A fortiori, $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq(\Gamma,\widetilde{v})$, whence $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subset(\Gamma^{B},\widetilde{v})$. Now in $(\Gamma^{B},\widetilde{v})$, we can read $g$ on a walk $v^{\prime}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$s$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime\prime}$ such that $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$s$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime\prime}$ is a subwalk of $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$ and where $v^{\prime}\neq v^{\prime\prime}$, otherwise determinism would imply $\widetilde{v}\cdot\mathring{g}=\widetilde{v}$. In particular, $g$ is non trivial in $G(\Theta_{B})$.
The previous argument may easily be adapted when considering a finite family of disjoint complete context-free inverse graphs. ∎
Note that the previous proposition gives a constructive way to compute context-free inverse graphs defining a finitely generated subgroup of a transition group of a context-free graphs $\Gamma$.
Proposition 4.15.
The class of CF-TR is closed by taking direct products.
Proof.
Let $\mathcal{G}(\Gamma_{i})$ be a transition group of a context-free inverse $A_{i}$-digraph with $i=1,2$. We can assume that $A_{1}\cap A_{2}$ is empty, otherwise we change accordingly the labels in $A_{2}$. Now we consider $\widehat{\Gamma}_{1}\sqcup\widehat{\Gamma}_{2}$ where $\widehat{\Gamma}_{i}$ is a $A_{1}\cup A_{2}$-digraph in which for every vertex there are $|A_{j}|$ loops each labeled by an element in $A_{j}$ with $j\neq i$. It is straightforward that $\mathcal{G}(\widehat{\Gamma}_{1}\sqcup\widehat{\Gamma}_{2})$ is the desired direct product. ∎
Definition 4.16.
Let $H$ be a finitely generated group which is a finite index subgroup of $G$ and let $T$ be a right transversal for $G$ with respect to $H$ containing the identity. For each $y\in A\cup T-\{\mathds{1}_{G}\}$ and $t\in T$, we fix a word $h_{t,y}\in A^{*}$ such that $ty=_{G}h_{t,y}\widetilde{t}$ for some $\widetilde{t}\in T$. We define the Schreier automaton $\mathcal{A}$ over $A\cup T$ to be the finite state automaton with set of states $T$ , initial state $\mathds{1}_{G}\in T$ and $t\smash{\stackrel{{\scriptstyle\text{\tiny{$y\mid h_{t,y}$}}}}{{\relbar\joinrel\longrightarrow}}}\widetilde{t}$ when $ty=_{G}h_{t,y}\widetilde{t}$.
$t\smash{\stackrel{{\scriptstyle\text{\tiny{$y\mid h_{t,y}$}}}}{{\relbar\joinrel\longrightarrow}}}\widetilde{t}\ \Longleftrightarrow\widetilde{t}\ \smash{\stackrel{{\scriptstyle\text{\tiny{$y^{-1}\mid h^{-1}_{t,y}$}}}}{{\relbar\joinrel\longrightarrow}}}t,$
without changing the groups involved $H$ and $G$. The way we do so is by enlarging the alphabet for the labels from $A\cup T$ to $A\cup\{T,T^{-1}\}$ (remember $A$ is already inverse by assumption) and by adding the inverse transitions $\widetilde{t}\ \smash{\stackrel{{\scriptstyle\text{\tiny{$y^{-1}\mid h^{-1}_{t,y}$}}}}{{\relbar\joinrel\longrightarrow}}}t$. The fact that we are still dealing with $H$ and $G$ follows from the equation $\widetilde{t}y^{-1}=_{G}h^{-1}_{t,y}t$, which is true if and only if $ty=_{G}h_{t,y}\widetilde{t}$. We call it the inverse Schreier automaton and with a slight abuse of notation we denote it again by $\mathcal{A}$.
Definition 4.17.
Let $\Gamma$ be a complete inverse $A$-digraph and let $H$ be the associated transition group. Consider $G$ to be a group with $H$ as a finite index subgroup and $\mathcal{A}$ the corresponding Schreier automaton over $A\cup T$. The inflated graph $\mathcal{A}\ltimes(\Gamma,x)$ with $x\in V(\Gamma)$ is the rooted $A\cup T$-digraph defined as follows: the set of vertices is $T\times V(\Gamma)$, a generic edge is $(t,p)\smash{\stackrel{{\scriptstyle\text{\tiny{$y$}}}}{{\relbar\joinrel\longrightarrow}}}(t^{\prime},p^{\prime})$ with $t\smash{\stackrel{{\scriptstyle\text{\tiny{$y\mid u$}}}}{{\relbar\joinrel\longrightarrow}}}t^{\prime}\in\mathcal{A}$ and $p\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}$ a walk in $\Gamma$ and the root is $(\mathds{1}_{G},x)$.
It is straightforward that $\mathcal{A}\ltimes\Gamma$ is complete and if $\Gamma$ is context-free, so is $\mathcal{A}\ltimes\Gamma$ (see the proof of [24, Lemma 11]).
Proposition 4.18.
The class of CF-TR is closed by taking finite index overgroups.
Proof.
Let $\{\Delta_{1},\ldots,\Delta_{N}\}$ be the frontier points, and arguing as we did in Proposition 4.14 with a slight abuse of notation, we identify the frontier vertices of $\Delta_{i}$ with the frontier vertices of an end cone $\Gamma(v,x_{0})$ of $\Gamma$ that is end-isomorphic to $\Gamma_{i}$. We claim that $G$ is the transition group of
  1. (1)
  2. (2)
    for all $t\in T$ there exists $h\in H$ s.t. $g=_{G}ht$ and $h$ is non trivial in $H$.
Suppose there exists $i$ and $v\in\Delta_{i}$ as in the hypothesis, then $(t,p)$ with $t\in T$ and $p\in V(\Gamma)$ is such that $(t,p)\cdot g=(t^{\prime},p^{\prime})$ and in $\mathcal{A}$ we have $t\smash{\stackrel{{\scriptstyle\text{\tiny{$g\mid h$}}}}{{\relbar\joinrel\longrightarrow}}}t^{\prime}$. If $t\neq t^{\prime}$, then in $\mathcal{A}$ we have $t\smash{\stackrel{{\scriptstyle\text{\tiny{$g\mid h$}}}}{{\relbar\joinrel\longrightarrow}}}t^{\prime}\neq t$ hence we have $(t,p)\cdot g=(t^{\prime},p\cdot h)\neq(t,p)$ in $\mathcal{A}\rtimes(\Gamma,v)$ for some $i$ and $v\in\Delta_{i}$. Otherwise by definition we have $p\smash{\stackrel{{\scriptstyle\text{\tiny{$h$}}}}{{\relbar\joinrel\longrightarrow}}}p\cdot h\neq p$ in $\Gamma$ and $h$ is non trivial in $H$. So, by Lemma 4.5, there exists $\widetilde{p}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{h}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(\widetilde{p},x_{0})$ with $\mathring{h}\in S_{A}(h)$. Therefore there exists $p^{\prime}\in V(\Gamma,v)$ for some $i$ and $v\in\Delta_{i}$ such that $p^{\prime}\cdot h\neq p^{\prime}$ and $p^{\prime}\smash{\stackrel{{\scriptstyle\text{\tiny{$h$}}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}\cdot h\subseteq\Gamma(v,x_{0})$. Finally, in $\mathcal{A}\rtimes(\Gamma,v)$ we have $(\mathds{1}_{G},v)\cdot u=(t,p^{\prime})$ where $u$ is such that there exists $\mathds{1}_{G}\smash{\stackrel{{\scriptstyle\text{\tiny{$u\mid z$}}}}{{\relbar\joinrel\longrightarrow}}}t$ in $\mathcal{A}$ with $v\smash{\stackrel{{\scriptstyle\text{\tiny{$z$}}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}\subseteq(\Gamma,v)$ and $(t,p^{\prime})\cdot g=(t,p^{\prime}\cdot h)\neq(t,p^{\prime})$ since by hyphotesis $tg=_{G}ht$ and $p^{\prime}\cdot h\neq p^{\prime}$. ∎
Remark 4.19.
In particular, this means that any finitely generated subgroup of a virtual direct product of free groups is CF-TR.
@@ 4.1. An example
As a concrete example of a transition group of an inverse graph we consider the inverse $A$-digraph $\Omega$ on the alphabet $A=\{a,a^{-1},b,b^{-1},c,c^{-1}\}$ (see Fig. 1). The set of vertices $V(\Omega)=\{p_{i},q_{i},i\in\mathbb{Z}\}$ and the set $E(\Omega)$ of edges is given by:
$p_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}p_{i+1},\;p_{i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}p_{i},\,q_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b$}}}}{{\relbar\joinrel\longrightarrow}}}q_{i+1},\;q_{i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$b^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}q_{i},\quad\forall i\in\mathbb{Z}$
$p_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b$}}}}{{\relbar\joinrel\longrightarrow}}}p_{i},\;p_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}p_{i},\,q_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}q_{i},\;q_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}q_{i},p_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}q_{i},q_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}p_{i},x\in\{c,c^{-1}\}\quad\forall i\in\mathbb{Z}$
@@
Figure 1. The inverse graph $\Omega$. The loops in red are labelled by $a,a^{-1}$, in blue by $b,b^{-1}$.
Note that $\Omega$ is clearly quasi-transitive and it is quasi-isometric to a tree, thus by [24, Theorem 2] it is a context-free graph. The next proposition characterizes the transition group defined by $\Omega$.

### v3

@@ 4.1. Transition groups of context-free graphs
Let us prove the characterization. As observed in Example 4.3, it is straightforward to check that the transition group of the Cayley graph $\operatorname{Cay}(G;A)$ is isomorphic to $G$. Thus, if there is a covering $\psi:\operatorname{Cay}(G;A)\to\Gamma$ onto a context-free graph $\Gamma$, then by the first part of this proposition $\mathcal{G}(\Gamma)\simeq G$. Conversely, if $G=\mathcal{G}(\Gamma)$ and we fix a root $x_{0}$ in $\Gamma$, then the morphism $\psi:\operatorname{Cay}(G;A)\to\Gamma$ defined on the vertices by $\psi(v)=y$ whenever $1\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g$}}}{{\relbar\joinrel\longrightarrow}}}v$ is a walk in $\operatorname{Cay}(G;A)$ and $y=x_{0}\cdot g$ and extended to a morphism of $A$-graphs, is a covering that is also transition-preserving. ∎
@@ 4.2. An example
As a concrete example of a transition group of a context-free graph we consider the $A$-graph $\Omega$ on the alphabet $A=\{a,a^{-1},b,b^{-1},c,c^{-1}\}$ (see Fig. 1). The set of vertices $V(\Omega)=\{p_{i},q_{i},i\in\mathbb{Z}\}$ and the set $E(\Omega)$ of edges is given by:
$p_{i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}p_{i+1},\;p_{i+1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a^{-1}$}}}{{\relbar\joinrel\longrightarrow}}}p_{i},\,q_{i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$b$}}}{{\relbar\joinrel\longrightarrow}}}q_{i+1},\;q_{i+1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$b^{-1}$}}}{{\relbar\joinrel\longrightarrow}}}q_{i},\quad\forall i\in\mathbb{Z}$
$p_{i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$b$}}}{{\relbar\joinrel\longrightarrow}}}p_{i},\;p_{i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$b^{-1}$}}}{{\relbar\joinrel\longrightarrow}}}p_{i},\,q_{i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}q_{i},\;q_{i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a^{-1}$}}}{{\relbar\joinrel\longrightarrow}}}q_{i},p_{i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$x$}}}{{\relbar\joinrel\longrightarrow}}}q_{i},q_{i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$x$}}}{{\relbar\joinrel\longrightarrow}}}p_{i},x\in\{c,c^{-1}\}\quad\forall i\in\mathbb{Z}$
@@
Figure 1. The graph $\Omega$. The loops in red are labeled by $a,a^{-1}$, in blue by $b,b^{-1}$.
Note that $\Omega$ is clearly quasi-transitive and it is quasi-isometric to a tree, thus by [27, Theorem 2] it is a context-free graph. The next proposition characterizes the transition group defined by $\Omega$.
@@
4.3. Closure properties
In this section we show that this class is closed under taking finitely generated subgroups, direct products, and finite extensions. We fix $H$ a subgroup of $G$ that is finitely generated by $B\subseteq A^{*}$ and we set $n:=\max_{b\in B}\ell(b)$ with $\ell(b)$ the length in terms of $A$.
Lemma 4.16.
Let $g\in H$ be a nontrivial element. Then, there exist an end-cone $\Gamma(v,x_{0})$ and a vertex $\widetilde{v}\in\Gamma(v,x_{0})$ with $d(v,\widetilde{v})\leq n$ such that the infinite walk $\widetilde{v}\xrightarrow{\mathring{g}^{\omega}}\ldots\subseteq\Gamma(v,x_{0})$ for some $\mathring{g}\in S_{B}(\{g\})$ and $\widetilde{v}\cdot\mathring{g}\neq\widetilde{v}$.
Proof.
By Lemma 4.7, there exists an end-cone $\Gamma(v,x_{0})$ and a cyclic shift $\widehat{g}\in S(\{g\})$ such that the infinite walk $\pi=x\xrightarrow{\widehat{g}^{\omega}}\ldots$ is contained in $\Gamma(v,x_{0})$. By the definition of $n$, possibly removing a prefix of $\pi$ of length at most $n$, we may choose a vertex $\widetilde{v}\in\pi$ such that $d(v,\widetilde{v})\leq n$, colored by a prefix $g^{\prime}$ of $g$, with $g=g^{\prime}g^{\prime\prime}$, and set $\mathring{g}=g^{\prime\prime}g^{\prime}\in S_{B}(\{g\})$. By construction, $\mathring{g}$ satisfies the required properties, namely $\widetilde{v}\xrightarrow{\mathring{g}^{\omega}}\ldots\subseteq\Gamma(v,x_{0})$ and $\widetilde{v}\cdot\mathring{g}\neq\widetilde{v}$. ∎
Proposition 4.17.
The class of CF-TR is closed by taking finitely generated subgroups.
Proof.
Suppose first that we are dealing with one complete context-free graph, so that $H$ is a finitely generated subgroup of $\mathcal{G}(\Gamma)$, where $\Gamma$ is a graph that is also context-free. Let $B\subseteq A^{*}$ be the involutive closed generating set and let $\Delta_{1},\ldots,\Delta_{N}$ be the frontier vertices of the end-cone types $\{\Gamma_{1},\ldots,\Gamma_{N}\}$. We set $\Delta_{i}^{n}:=\{v\in\Gamma_{i}\mid d(x,v)\leq n\mbox{ for some }x\in\Delta_{i}\}$. With a slight abuse of notation we identify the frontier vertices of $\Delta_{i}$ with the frontier vertices of an end cone $\Gamma(v,x_{0})$ of $\Gamma$ that is end-isomorphic to $\Gamma_{i}$. We do the same for the set $\Delta_{i}^{n}$. By definition of transition group, we need to show that $H$ is the transition group of a finite union of complete context-free $B$-graphs $\Theta_{B}$. We claim
$\Theta_{B}=\bigcup_{i=1}^{N}\bigcup_{v\in\Delta_{i}^{n}}(\Gamma^{B},v)$
where $(\Gamma^{B},v)$ is given by the vertices $y$ of $\Gamma$ such that there exists $w\in B^{*}$ and $v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w$}}}{{\relbar\joinrel\longrightarrow}}}y$. The edges are $y_{1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w$}}}{{\relbar\joinrel\longrightarrow}}}y_{2}$ whenever $w\in B$ and there exists a walk labeled by $w$ from $y_{1}$ to $y_{2}$ in $\Gamma$. It is not difficult to see that since $B$ is involutive closed and $\Gamma$ is inverse, then also $(\Gamma^{B},v)$ is inverse. Moreover, it is complete and context-free.
Let us show that the complement of the word problem of $H$ and $\mathcal{G}(\Theta_{B})$ coincide, as this will of course imply that their word problems coincide. Let $g\in G(\Theta_{B})$ be a non trivial element, then there exists $\widetilde{v}\in(\Gamma^{B},v)$ for some $v\in\Delta_{i}^{n}$ such that $\widetilde{v}\cdot g\neq\widetilde{v}$. It follows that $\widetilde{v}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g$}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime}$ with $\widetilde{v}\neq v^{\prime}$ is also a walk in $\Gamma$ and so $g\neq\mathds{1}_{H}$ in $H$.
[…]
The previous argument may easily be adapted when considering a finite family of disjoint complete context-free graphs. ∎
Note that the previous proposition gives a constructive way to compute context-free graphs defining a finitely generated subgroup of a transition group of a context-free graphs $\Gamma$.
Proposition 4.18.
The class of CF-TR is closed by taking direct products.
Proof.
Let $\mathcal{G}(\Gamma_{i})$ be a transition group of a context-free $A_{i}$-graph with $i=1,2$. We can assume that $A_{1}\cap A_{2}$ is empty, otherwise we change accordingly the labels in $A_{2}$. Now we consider $\widehat{\Gamma}_{1}\sqcup\widehat{\Gamma}_{2}$ where $\widehat{\Gamma}_{i}$ is a $A_{1}\cup A_{2}$-graph in which for every vertex there are $|A_{j}|$ loops each labeled by an element in $A_{j}$ with $j\neq i$. It is straightforward that $\mathcal{G}(\widehat{\Gamma}_{1}\sqcup\widehat{\Gamma}_{2})$ is the desired direct product. ∎
Definition 4.19.
Let $H$ be a finitely generated group which is a finite index subgroup of $G$ and let $T$ be a right transversal for $G$ with respect to $H$ containing the identity. For each $y\in A\cup T-\{\mathds{1}_{G}\}$ and $t\in T$, we fix a word $h_{t,y}\in A^{*}$ such that $ty=_{G}h_{t,y}\widetilde{t}$ for some $\widetilde{t}\in T$. We define the Schreier automaton $\mathcal{A}$ over $A\cup T$ to be the finite state automaton with set of states $T$ , initial state $\mathds{1}_{G}\in T$ and $t\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$y\mid h_{t,y}$}}}{{\relbar\joinrel\longrightarrow}}}\widetilde{t}$ when $ty=_{G}h_{t,y}\widetilde{t}$.
$t\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$y\mid h_{t,y}$}}}{{\relbar\joinrel\longrightarrow}}}\widetilde{t}\ \Longleftrightarrow\widetilde{t}\ \smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$y^{-1}\mid h^{-1}_{t,y}$}}}{{\relbar\joinrel\longrightarrow}}}t,$
without changing the groups involved $H$ and $G$. The way we do so is by enlarging the alphabet for the labels from $A\cup T$ to $A\cup\{T,T^{-1}\}$ (remember $A$ is already inverse by assumption) and by adding the inverse transitions $\widetilde{t}\ \smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$y^{-1}\mid h^{-1}_{t,y}$}}}{{\relbar\joinrel\longrightarrow}}}t$. The fact that we are still dealing with $H$ and $G$ follows from the equation $\widetilde{t}y^{-1}=_{G}h^{-1}_{t,y}t$, which is true if and only if $ty=_{G}h_{t,y}\widetilde{t}$. We call it the inverse Schreier automaton and with a slight abuse of notation we denote it again by $\mathcal{A}$.
Definition 4.20.
Let $\Gamma$ be a complete $A$-graph and let $H=\mathcal{G}(\Gamma)$ be the associated transition group. Consider $G$ to be a group with $H$ as a finite index subgroup and $\mathcal{A}$ the corresponding Schreier automaton over $A\cup T$. The inflated graph $\mathcal{A}\ltimes(\Gamma,x)$ with $x\in V(\Gamma)$ is the rooted $A\cup T$-graph defined as follows: the set of vertices is $T\times V(\Gamma)$, a generic edge is $(t,p)\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$y$}}}{{\relbar\joinrel\longrightarrow}}}(t^{\prime},p^{\prime})$ with $t\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$y\mid u$}}}{{\relbar\joinrel\longrightarrow}}}t^{\prime}\in\mathcal{A}$ and $p\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}$ a walk in $\Gamma$ and the root is $(\mathds{1}_{G},x)$.
Intuitively, the inflated graph simulates a cascade composition of the Schreier automaton $\mathcal{A}$ with the graph $\Gamma$: each transition of $\mathcal{A}$ labelled by $y\mid u$ triggers the traversal of the word $u$ inside $\Gamma$. This construction is not genuinely new, it can be viewed as a composition of sequential machines (transducers) in disguise, where the graph $\Gamma$ itself is interpreted as a transducer. It is straightforward to see that $\mathcal{A}\ltimes\Gamma$ is inverse and if $\Gamma$ is context-free, so is $\mathcal{A}\ltimes\Gamma$ (see the proof of [27, Lemma 11]).
Proposition 4.21.
The class of CF-TR is closed by taking finite index overgroups.
Proof.
Let $\{\Delta_{1},\ldots,\Delta_{N}\}$ be the frontier vertices, and arguing as we did in Proposition 4.17 with a slight abuse of notation, we identify the frontier vertices of $\Delta_{i}$ with the frontier vertices of an end cone $\Gamma(v,x_{0})$ of $\Gamma$ that is end-isomorphic to $\Gamma_{i}$. We claim that $G$ is the transition group of
  1. (1)
  2. (2)
    for all $t\in T$ there exists $h\in H$ s.t. $g=_{G}ht$ and $h$ is non trivial in $H$.
Suppose there exists $i$ and $v\in\Delta_{i}$ as in the hypothesis, then $(t,p)$ with $t\in T$ and $p\in V(\Gamma)$ is such that $(t,p)\cdot g=(t^{\prime},p^{\prime})$ and in $\mathcal{A}$ we have $t\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g\mid h$}}}{{\relbar\joinrel\longrightarrow}}}t^{\prime}$. If $t\neq t^{\prime}$, then in $\mathcal{A}$ we have $t\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g\mid h$}}}{{\relbar\joinrel\longrightarrow}}}t^{\prime}\neq t$ hence we have $(t,p)\cdot g=(t^{\prime},p\cdot h)\neq(t,p)$ in $\mathcal{A}\rtimes(\Gamma,v)$ for some $i$ and $v\in\Delta_{i}$. Otherwise by definition we have $p\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$h$}}}{{\relbar\joinrel\longrightarrow}}}p\cdot h\neq p$ in $\Gamma$ and $h$ is non trivial in $H$. So, by Lemma 4.7, there exists $\widetilde{p}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$\mathring{h}^{\omega}$}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(\widetilde{p},x_{0})$ with $\mathring{h}\in S_{A}(h)$. Therefore there exists $p^{\prime}\in V(\Gamma,v)$ for some $i$ and $v\in\Delta_{i}$ such that $p^{\prime}\cdot h\neq p^{\prime}$ and $p^{\prime}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$h$}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}\cdot h\subseteq\Gamma(v,x_{0})$. Finally, in $\mathcal{A}\rtimes(\Gamma,v)$ we have $(\mathds{1}_{G},v)\cdot u=(t,p^{\prime})$ where $u$ is such that there exists $\mathds{1}_{G}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u\mid z$}}}{{\relbar\joinrel\longrightarrow}}}t$ in $\mathcal{A}$ with $v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$z$}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}\subseteq(\Gamma,v)$ and $(t,p^{\prime})\cdot g=(t,p^{\prime}\cdot h)\neq(t,p^{\prime})$ since by hyphotesis $tg=_{G}ht$ and $p^{\prime}\cdot h\neq p^{\prime}$. ∎
Remark 4.22.
This means that any finitely generated subgroup of a virtual direct product of free groups is CF-TR. In particular, finitely generated virtually abelian groups are in CF-TR.